### v1

# The Detection of a Possible Exoplanet Orbiting KIC 1718360 Using Machine Learning

Abstract
This paper presents the detection of a periodic dimming event in the lightcurve of the G1.5IV-V type star KIC 1718360. This is based on visible-light observations conducted by both the TESS and Kepler space telescopes. Analysis of the data points toward a possible orbiting body with a radius of approximately 1.048 $R_{\oplus}$ with a period of 2.938 days, as well as a semi-major axis of 0.04 AU. The initial observation was made in Kepler Quarter 16 data using the One-Class SVM machine learning method. Subsequent observations by the TESS space telescope corroborate these findings. While still requiring further data to validate, these results may contribute to a growing body of data of Earthlike planets with short-period orbits.
Introduction
At the time of writing, over 5,200 exoplanets have been confirmed11footnotemark: 1, and thousands more are candidates. This amounts to a wealth of data on many different planet types, including ones not found in our solar system. However, in discovering these planets, there is still a prevalence of observing lightcurves manually. This is often a painstaking process, and one that can be automated using machine learning techniques.
This paper presents the finding of a periodic dip in the lightcurve of the star KIC 1718360 that was detected in both Kepler and TESS data using a One-Class Support Vector Machine (SVM) model. This observation is indicative of an exoplanet companion orbiting the G1.5IV-V type star. The star is thought to have a size comparable to that of the sun, with an estimated size of 0.91 $R_{\odot}$
Procedure
Given the fact that there are expected to be a high number of exoplanets in the universe22footnotemark: 2, and the chances being relatively high that some transit signals might be overlooked if lightcurves were to be labeled as either having a transit or not33footnotemark: 3, traditional machine learning methods could not be effectively used. These methods require a positive dataset, which in this case would be lightcurves that do contain exoplanets, and a negative dataset, which would be lightcurves that do not contain exoplanets. For the aforementioned reasons, instead of training a model on data labeled as positive or negative, alternative techniques had to be used.
The technique used in this paper revolves around anomaly detection algorithms, namely the One-Class SVM model. In the following procedure, approximately 1,500 datafiles were provided in the form of Kepler Object of Interest (KOI) lightcurves. To begin the process of training the model, the the Pre-Search Data Conditioning Simple Aperture Photometry (PDCSAP) flux data provided by the mission was processed into a 2-dimensional array. This array contained the raw flux data for each target star in a 2-dimensional format, with each target having a separate row. This was done to put the data in a format suitable for machine learning using the selected algorithm. This data was then normalized and detrended using a Savitzky-Golay filter44footnotemark: 4. The machine learning model used was trained to find lightcurves that have high similarity to the KOI lightcurve data. The algorithm’s similarity parameters included a periodic dip in the lightcurve as well as a dip significant enough to be a planet.
After a period of validation and parameter adjustment, the model was deployed onto the Kepler Quarter 16 data. The star KIC 1718360, among others, was flagged as being highly similar to KOI lightcurves, despite not being on the KOI catalog themselves55footnotemark: 5. Upon manual inspection, KIC 1718360 was determined to be the most probable exoplanet host, since it has the most obvious transit signal as well as a signal that clearly recurs in all Kepler data. Expected orbital parameters of the exoplanet were calculated from the lightcurve data, including its expected diameter and its orbital distance.
Upon closer inspection, the possible transit signal was found to be present in all of Kepler’s observations, ranging from Kepler’s first light in 2009 to the end of the Kepler mission four years later. Evidence corroborating this finding was sourced from the Transiting Exoplanet Survey Satellite (TESS). The TESS observations referenced were conducted in 2022, with virtually exactly the same period and expected parameters. In total, this amounts to a thirteen year period in which the same signal was detected around this star. Given that there are no catalogued exoplanets orbiting KIC 1718360, and it is not classified as a variable, this star is a great exoplanet host candidate.
Algorithms
A One-Class Support Vector Machine (SVM) was used to make predictions on the Kepler dataset. One-Class SVMs are primarily anomaly-detection algorithms, and have applications in finding outlying datapoints in much larger datasets66footnotemark: 6. In this scenario, every lightcurve was fed into the model and treated as an individual datapoint. Outlying datapoints were treated as possible transit signals and further analyzed, while the non-outlying lightcurves were assumed to contain no transits.
The model used a nonlinear, polynomial kernel with a degree of 4 to run predictions. When predictions were run, the majority of lightcurves were found to be non-outliers, with a significant minority (on average around 6%) of lightcurves were found to be anomalous. This method proved effective in finding periodic dips in lightcurves; a large number of variable stars were flagged by the model.
Data Analysis
Figure 1: This is the original lightcurve detected by the model from Kepler’s Quarter 16 observation. Quarter 16 lasted approximately 80 days.
Figure 1 depicts the lightcurve that the model flagged as possibly being a transit. While some long-term variation in the brightness is observed, it does not seem to impact the recurrence of a signal with a period of 2.9376 days. While the lightcurve shares some characteristics of variable star lightcurves, this star is not catalogued as a variable77footnotemark: 7. While it is possible that it is an uncatalogued variable, the star’s relative proximity to Earth (around 978 ly), and the extent to which we’ve catalogued nearby variables88footnotemark: 8makes this unlikely.
Figure 2: This is the lightcurve in Figure 1 folded and scattered by a period of approximately 2.94 days. It was folded 26 times.
Figure 2 shows the previous lightcurve folded and scattered by the suspected period. This was important in calculating the expected planetary parameters. It is highly unlikely that this signal is caused by a starspot99footnotemark: 9, as it appears not only in Kepler’s 4-year observation period, but also in 2022 as well.
Figure 3: A Box Least Squares (BLS) periodogram created based on the Figure 1 data. Note the peak near 2.94 days, signifying a recurring signal of that period.
The Box Least Squares (BLS) periodogram in Figure 3 shows a clear peak at approximately 2.9 days. This periodogram was used to find the period of the suspected transit.
Figure 4: TESS observations of KIC 1718360. Note the same trend observed in the Kepler data is also seen here.
When the TESS catalog was queried with the star’s TIC identifier (TIC 122673562), the lightcurve in Figure 4 was produced. The lightcurve was produced from TESS Sector 55 observations, which took place from August 5 to September 1, 2022. The fact that exactly the same transit signal has persisted from 2009 to 2022 strongly points toward a possible planetary nature for the signal. The signal remaining unchanged for a 13 year period virtually rules out using starspots or other short-term solar phenomena to explain the signal.
Figure 5: The lightcurve in Figure 4 folded and scattered by a period of approximately 2.94 days. Note the high similarity, yet considerably less noise, compared to the data in Figure 2.
Figure 6: The lightcurve of Kepler-8b folded and scattered by its period of 3.52 days. Note the similarity in the shape of the dip to the lightcurve in Figure 5.
Figure 5 depicts the TESS Sector 15 data folded and scattered by the suspected period. The resulting lightcurve is highly similar to the lightcurve in Figure 2, yet it has much less instrumental noise than the original observation. Whereas the graph in Figure 2 has noise along the depth of the lightcurve, Figure 5 strongly resembles a classical transit signal, like the one in Figure 6.
Parameter Analysis
The following are the planetary parameters calculated from the lightcurve in Figure 2.
Period Radius Semi-Major Axis Volume
2.9376 days 1.0489 $R_{\oplus}$ 0.04 AU 1.254E+21 $m^{3}$
Table 1: The expected planetary parameters for the suspected transit.
If the signal is indeed caused by an exoplanet, it is a planet very nearly the size of Earth orbiting a star highly similar to the sun. The only difference between the Earth-Sun system and the suspected exoplanet, naturally, are the extremely high temperatures to be expected on the planet’s surface.
The planet’s relatively small radius is indicative of its suspected rocky composition, to date no Earth-sized gas giants have been detected1010footnotemark: 10. Making a few broad assumptions about the albedo of the planet (10%, consistent with the idea that the planet has no atmosphere and is made primarily of rock), an average surface temperature of approximately 1330 Kelvin can be calculated. At this temperature, lead would melt.
While this planet is in extreme proximity to its host star, this is fully possible given the latest exoplanet research and it is not without precedent1111footnotemark: 11. If this signal is an exoplanet, its existence is fully plausible and within the range of exoplanet types already discovered 1212footnotemark: 12.
Conclusions
This research provides compelling evidence for the presence of a possible exoplanet around KIC 1718360. Analysis of the periodic dimming event in the lightcurve, observed by both the TESS and Kepler space telescopes, suggests the possibility of a small planet with a radius of approximately 1.049 $R_{\oplus}$ and a period of 2.938 days orbiting the target star.
While it is not possible to definitively conclude the nature of the signal from lightcurve data alone, the data collected strongly suggests the possibility of an exoplanet. The fact that multiple space telescopes across long periods of time corroborate the finding makes the signal unlikely to be any kind of instrumental noise or short term stellar fluctuations.
This find requires future study to validate. Until then, KIC 1718360 remains a possible exoplanet.